# AIGC for Industrial Time Series: From Deep Generative Models to Large Generative Models

### v1

@@ Abstract
With the remarkable success of generative models like ChatGPT, Artificial Intelligence Generated Content (AIGC) is undergoing explosive development. Not limited to text and images, generative models can generate industrial time series data, addressing challenges such as the difficulty of data collection and data annotation. Due to their outstanding generation ability, they have been widely used in Internet of Things, metaverse, and cyber-physical-social systems to enhance the efficiency of industrial production. In this paper, we present a comprehensive overview of generative models for industrial time series from deep generative models (DGMs) to large generative models (LGMs). First, a DGM-based AIGC framework is proposed for industrial time series generation. Within this framework, we survey advanced industrial DGMs and present a multi-perspective categorization. Furthermore, we systematically analyze the critical technologies required to construct industrial LGMs from four aspects: large-scale industrial dataset, LGMs architecture for complex industrial characteristics, self-supervised training for industrial time series, and fine-tuning of industrial downstream tasks. Finally, we conclude the challenges and future directions to enable the development of generative models in industry.
@@ I Introduction
Since generative models have been applied in various fields and have shown promising results in industrial time series applications, there have been recent studies analyzing and surveying DGMs. Comprehensive reviews, such as the survey[42], provide profound insights into diffusion models across domains such as image synthesis, time series prediction, and natural language generation. Additionally, domain-specific reviews like [43] presented a review about the application of DGMs in the industry. Similarly, with the rapid development of diffusion models, there are some reviews on diffusion models in specific fields such as biological signals[44] or computer vision[45]. Due to the significance of time series in industry, literature [3] has surveyed time series data modeling issues. In addition to DGMs with specific-domain, specific-task, and specified generation ability, LGMs with multi-domain, multi-task, and generalized generation ability have made significant progress, as shown in Fig. 1. Some recent studies[46, 47] have investigated the LGMs in general applications.
@@
3) This paper proposes a systematic analysis of the critical technologies required to construct industrial LGMs that are suitable for industrial time series data.
The organization of the rest of the sections is as follows. In Section II, the commonly used DGMs are introduced. In Section III, the definition, characteristics, and requirements of industrial time series generation are presented. In Section IV, the proposed DGM-based AIGC framework and the current state-of-the-art DGMs are introduced. In Section V, we present the typical industrial applications of DGMs. In Section VI, we analyze how to construct LGMs for industrial time series. Finally, Section VII concludes the challenges and future directions of generative models.
@@ II-A Autoregressive Model
An autoregressive model (AR)[50] is a particular regression model, wherein a value from the time series is subjected to a regression on previous values. It predicts future values in a sequence based on previous values. The autoregressive model of order $p$ (often termed $AR(p)$ model) is defined as:
@@
where $X_{t}$ is the value of the series at time $t$, $c$ is a constant, $\phi_{1},\phi_{2},\ldots,\phi_{p}$ are the parameters of the model, $\epsilon_{t}$ is white noise. The likelihood for an $AR(p)$ model given data $\{x_{1},x_{2},\ldots,x_{T}\}$ is:
$L(\phi;x_{1},x_{2},\ldots,x_{T})=\prod_{t=p+1}^{T}f(x_{t}|x_{t-1},x_{t-2},\ldots,x_{t-p})$ (2)
where $f$ is the conditional density of $x_{t}$ given past values.
For many autoregressive models, especially when the noise $\epsilon_{t}$ is Gaussian, the negative log-likelihood serves as loss function:
@@
Minimizing this loss function gives the maximum likelihood estimates of the parameters.
Autoregressive models are commonly used in industrial forecasting[51][52]. For example, Wang et al.[53] proposed a model, combining non-linear autoregressive neural network and empirical mode decomposition to address the problem of nonlinear time series analysis for predicting industrial iron ore import volumes.
@@ II-B Variational Autoencoder (VAE)
VAE is a generative model that extends the concept of the traditional auto-encoder (AE)[54] by introducing a probabilistic approach to describe the data generation process. It operates by first encoding an input into a distribution over a latent space and then sampling from this distribution to generate new data. Unlike a standard AE that simply compresses and reconstructs data, VAEs[16] are designed to produce novel samples. It is achieved by constructing a loss function that not only minimizes reconstruction error but also includes a term that regularizes the latent variable distribution, often encouraging it to approximate a standard normal distribution.
Given $x$ as the observed data, $z$ as the latent variable, and the following definitions:
  • $p(z)$: Prior distribution of the latent variable, typically chosen as a standard normal distribution $p(z)=N(z;0,I)$.
  • $q_{\phi}(z|x)$: Output of the encoder, i.e., the posterior distribution of the latent variable given the data with the weight $\phi$ of the encoder $\mathcal{E}$.
  • $p_{\theta}(x|z)$: Data generative model given the latent variable with the weight $\theta$ of the decoder $D$.
The VAE loss is defined as the sum of the reconstruction loss and the regularization loss:
$\mathcal{L}_{\text{VAE}}=\text{Reconstruction loss}+\text{Regularization loss}$ (4)
More specifically, the VAE loss is given by:
@@
VAE, like other deep learning techniques, has received a lot of attention from academia since its inception, but its application in industry may take some time as industrial applications usually require the maturity and stability of the technology. VAE has been successively applied to the industry for some anomaly detection[55][56][57], etc.
@@ II-C Generative Adversarial Network (GAN)
The Generative Adversarial Networks (GANs) were first proposed by Goodfellow et al.[15] in 2014. It is an unsupervised learning algorithm that involves the interaction of two independent neural networks against each other to achieve learning, which can be characterized as a zero-sum game. GANs do not require prior knowledge and do not require assumptions that the sample obeys a certain distribution.
A typical generative adversarial network consists of two parts: a discriminator $D$ and a generator $G$. The generator aims to generate synthetic data samples, whereas the discriminator strives to differentiate between actual and synthetic data. Hence, the objective function of GAN comprises maximizing the loss of the discriminator and minimizing the loss of the generator. Let $x$ represent real data, random noise $z$ is sampled from a prior distribution $p_{z}$, $G(z)$ is the synthetic sample generated by the generator from the noise $z$, $D(x)$ is the probability given by the discriminator that data $x$ is real. Then we have:
Generator Loss :
$L_{G}=-\mathbb{E}_{z\sim p_{z}(z)}[\log D(G(z))]$ (6)
Discriminator Loss:
$L_{D}=-\mathbb{E}_{x\sim p_{\text{data}}(x)}[\log D(x)]-\mathbb{E}_{z\sim p_{z}(z)}[\log(1-D(G(z)))]$ (7)
Hence, the objective function can be expressed as:
@@
In the evolution of GANs within industrial applications, key milestones are marked by their use in various domains over the years. In 2015, GANs were first leveraged for data augmentation purposes[58]. The scope of their application expanded in 2017 when they were utilized for fault detection[59]. A significant advancement came in 2018 when Zhao et al.[60] pioneered the application of GANs in defect detection, a concept further elaborated in [61]. This progression underscores the increasing versatility and significance of GANs in enhancing industrial processes through innovative AI-driven solutions.
@@ II-D Diffusion Model
Diffusion models, a group of generative models inspired by the physical process of diffusion, have become increasingly prominent in the field of machine learning recently. The Denoising Diffusion Probabilistic Model (DDPM)[19] is beginning to receive widespread academic attention. It is noteworthy to mention that before this study, Song Yang et al.[62] proposed a score-based diffusion model, and these two models were subsequently harmonized in the SDE study[63]. The core idea of Diffusion Models is to gradually transform the distribution of the data into a known simple distribution (e.g., a Gaussian distribution), and then sample from the simple distribution and reverse transform it back into the distribution of the data. And we implement both forward and backward processes through two Markov chains[64].
Given a distribution of data $p(x)$, a diffusion process can be defined at each time step $t$, and the data $x$ is perturbed to $x_{t}$. This process can be characterized as follows:
@@
where $\epsilon_{t}\sim\mathcal{N}(0,I)$ is a Gaussian noise and $\beta_{t}$ is a predefined noise factor.
A denoising function $q_{\theta}$ can be trained to reconstruct $x_{t-1}$ from $x_{t}$. Given a sample $x_{t}$ and its corresponding noise $\epsilon_{t}$, the output of the denoising function is a conditional distribution $q_{\theta}(x_{t-1}|x_{t})$.
The loss function of DDPM is based on the difference between the output of the denoising function and the true $x_{t-1}$. It can be expressed as:
@@
where the expectation is taken overall time steps and all possible $x_{t}$.
Diffusion models, initially emerging within computer vision[65], have expanded their utility to encompass sequential modeling[66, 67, 68], ventured into the audio domain[69], and bridged disciplines in AI for science[70]. These models have been adeptly integrated into various industrial sectors, tackling complex challenges such as the imputation of sensor data[71] and the nuanced task of anomaly detection[72].
@@ III-C Challenges in Industrial Time Series Acquisition
In summary, challenges associated with industrial time series generation include limited labeled time series, imbalanced industrial distribution, missing sensor time series values, industrial sensor noise, and industrial privacy concerns issues. Addressing these challenges is crucial for intelligent industrial applications.
@@ IV The proposed DGM-based AIGC framework and the current state-of-the-art DGMs
The DGM-baed AIGC framework is proposed for industrial data generation, as illustrated in Fig. 3. First, generative models learns the latent representations of industrial time series by modeling dynamic processes and correlations between variables, which have been introduced in Section II and Section III. Second, quality enhancement, sample augmentation, and privacy protection are performed to improve the quality and quantity of data, which will be discussed in Section IV. Finally, the generated samples will be applied to various industrial scenarios to address challenges such as sample scarcity in industrial scenarios, as presented in Section V. With the development of generative intelligence, there are many methods are proposed to tackle these issues. These research works can be mainly categorized into five: limited sample augmentation, imbalanced data synthesis, sensor signal imputation, sensor signal denoising, and privacy protection.
@@ IV-A Limited Sample Augmentation
Limited sample augmentation involves using sophisticated generative methods to enrich datasets where sample size is insufficient for effective model training. This approach is particularly crucial in scenarios where acquiring additional real-world data is challenging or impractical. By artificially generating realistic and diverse samples, this method enhances the dataset’s comprehensiveness, aiding in the development of more accurate and generalized machine learning models.
In recent industrial applications, the enhancement of model performance and data processing efficiency through limited sample augmentation has emerged as a significant trend. The bidirectional alignment network combined with VAE[73] effectively addresses the issue of scarce samples in fault diagnosis in thermal power plants, achieving generalization across rare and unseen fault categories. The DA-JITL framework[18] incorporates a causality-informed VAE, significantly enhancing the performance of JITL-based soft sensors. MSGAN[74] merges offline training with online real-time inference, employing an adaptive update strategy and gradient penalty in Wasserstein distance to generate high-quality false anomaly samples to compensate for the issue of data scarcity, thereby improving the accuracy of anomaly detection in industrial robotic sensors. Optimized diffusion models[24] and TimeDDPM[21] play crucial roles respectively in precise defect identification in CFRP structures and dynamic process soft sensor modeling by data augmentation, showcasing the versatility of these DGMs in limited sample augmentation.
@@ IV-B Imbalanced Data Synthesis
Imbalanced data synthesis refers to the process of artificially generating data to address the challenge of unequal distribution of classes within a dataset. This technique is pivotal in machine learning and data science, especially when dealing with scenarios where certain types of data (often representing minority classes) are scarce or underrepresented.
In the realm of industry, a myriad of innovative approaches are being explored to tackle the prevalent challenge of data imbalance in datasets. Among these, Fan et al.[35] introduces a groundbreaking Variational Autoencoder (VAE)-based imbalanced data synthesis method, tailored specifically for semiconductor fault detection. This method adeptly addresses the critical scarcity of defective wafer samples. Simultaneously, the DRL-GAN model[75] ingeniously amalgamates distributional reinforcement learning with GAN, thereby showcasing remarkable efficiency in generating balanced data distributions. Its application results in substantial improvements in various anomaly detection metrics, highlighting the model’s profound impact on the landscape of industrial anomaly detection. Furthermore, a Wasserstein conditional GAN[25] is proposed, which is enhanced with hierarchical feature matching. This novel approach is adept at tackling the long-tail distribution problem, a common quandary in the realm of bearing fault diagnosis. The implementation of this model demonstrates an effective and innovative method for addressing the data imbalance problem.
@@ IV-C Sensor Signal Imputation
The Vector Autoregressive Imputation Method (VAR-IM)[76] utilizes an integrated expectation-minimization and prediction error minimization algorithm, significantly improving imputation in multivariate time series, especially in scenarios involving missing signal data. Velasco-Gallego et al.[28] introduce a VAE-based sophisticated imputation method for sensor data from marine machinery. This approach yields a high determination coefficient, indicating its potential to improve the internet of ships and condition-based maintenance domains. Then in the realm of GANs, the Federated Transfer Missing Data Imputation method (FedTMI)[77] utilizes edge-computed GANs and federated transfer learning for heterogeneous missing data in industrial applications, while the Fine-Tuned Imputation GAN (FIGAN)[78] focuses on quality-related variables in soft sensors, interweaving improved data imputation with pseudo labeling. Other GAN models like the IM-GAN[79], SGAIN[80], Semi-GAN[81], ST-GAIN[82], and SGT-GAIN[83] also demonstrate superior capabilities in data imputation, each addressing specific challenges such as complex multivariate time series issues, structural health monitoring, semiconductor equipment fault detection, and advanced manufacturing data imputation. As for diffusion models, the DiffAD model[27] employs a denoising diffusion-based imputation method combined with a multi-scale state space model for time series anomaly detection. It effectively adapts to concentrated anomaly episodes and captures long-term dependencies. Similarly, the GD-GRU model[84] integrates a Gaussian diffusion process with a Gated Recurrent Unit for enhancing atmospheric environmental quality data repair.
@@ IV-D Sensor Signal Denoising
Starting with AR models, the studies[34] and [29] showcase diverse applications of AR in sensor signal denoising. Among these, Liu et al.[29] introduce an advanced approach for industrial robots, integrating wavelet regional correlation threshold denoising with autoregressive moving average models. This method significantly enhances signal clarity for acoustic emission data. In the domain of GAN models, a Bayesian nonparametric estimation within a CycleGAN architecture[30] is constructed for online sensor signal denoising in milling processes. Furthermore, the Att-DCDN is presented for seismic data denoising, combining an encoder-decoder structure within a GAN. This approach integrates synthetic and field seismic data, leveraging attribute-based constraints for noise attenuation. While Min et al. introduce a novel GAN-based approach[85] for seismic data denoising, combining a pre-trained deep denoising autoencoder with transfer learning to enhance training on limited field data.
@@
DiffAD.6[27] 2023 Anomaly Detection Diffusion Models
@@
Att-DCDN.[101] 2021 General Industrial Applications GAN

### v2

@@ Abstract
With the remarkable success of generative models like ChatGPT, Artificial Intelligence Generated Content (AIGC) is undergoing explosive development. Not limited to text and images, generative models can generate industrial time series data, addressing challenges such as the difficulty of data collection and data annotation. Due to their outstanding generation ability, they have been widely used in Internet of Things, metaverse, and cyber-physical-social systems to enhance the efficiency of industrial production. In this paper, we present a comprehensive overview of generative models for industrial time series from deep generative models (DGMs) to large generative models (LGMs). First, a DGM-based AIGC framework is proposed for industrial time series generation. Within this framework, we survey advanced industrial DGMs and present a multi-perspective categorization. Then, we systematically propose the roadmap to construct industrial LGMs from four aspects: large-scale industrial dataset, LGMs architecture for complex industrial characteristics, self-supervised training for industrial time series, and fine-tuning of industrial downstream tasks. Furthermore, we introduce an evaluation benchmark that systematically assesses fidelity, diversity, and utility. We include a case study on aircraft engine maintenance, demonstrating the application of DGMs in industrial predictive maintenance. Finally, we conclude the challenges and future directions to enable the development of generative models in industry.
@@ I Introduction
Since generative models have been applied in various fields and have shown promising results, there have been recent studies analyzing and surveying DGMs. Comprehensive reviews, such as the survey[42], provide profound insights into diffusion models across domains such as image synthesis, time series prediction, and natural language generation. Additionally, domain-specific reviews like [43] presented a review about the application of DGMs in the industry. Similarly, with the rapid development of diffusion models, there are some reviews on diffusion models in specific fields such as biological signals[44] or computer vision[45]. Due to the significance of time series in industry, literature [3] has surveyed time series data modeling issues. In addition to DGMs with specific-domain, specific-task, and specified generation ability, LGMs with multi-domain, multi-task, and generalized generation ability have made significant progress, as shown in Fig. 1. Some recent studies[46, 47] have investigated the LGMs in general applications.
@@
3) This paper proposes a systematic roadmap of the critical technologies required to construct industrial LGMs that are suitable for industrial time series data.
4) This paper establishes a comprehensive evaluation benchmark for generative models in industrial time series, systematically assessing fidelity, diversity, and utility. Additionally, a case study on aircraft engine maintenance is presented, demonstrating the effectiveness of DGMs in industrial predictive maintenance.
@@ II-A Autoregressive Model
An autoregressive model (AR) [50] predicts future values in a time series based on past observations. The autoregressive model of order $p$ (often termed $AR(p)$ model) is defined as:
@@
where $X_{t}$ is the value of the series at time $t$, $c$ is a constant, $\phi_{1},\phi_{2},\ldots,\phi_{p}$ are the model parameters, $\epsilon_{t}$ is white noise. For Gaussian noise, the negative log-likelihood serves as the loss function:
@@
Minimizing this function yields the maximum likelihood estimates of the parameters.
AR models are widely used in industrial forecasting. Wang et al. [51] introduced a hybrid model combining nonlinear autoregressive neural networks and empirical mode decomposition to improve industrial iron ore import volume predictions.
@@ II-B Variational Autoencoder (VAE)
A Variational Autoencoder (VAE) extends the traditional autoencoder (AE) by adopting a probabilistic approach to model data generation. Instead of directly encoding and reconstructing data, VAEs [16] map inputs to a latent distribution, from which new samples can be generated. The loss function consists of a reconstruction loss and a regularization term, encouraging the latent space to approximate a standard normal distribution.
Given observed data $x$ and latent variable $z$, the VAE loss is:
@@
Where $p(z)$ is the Prior distribution, typically $N(z;0,I)$, $q_{\phi}(z|x)$is the Encoder output, approximating the posterior distribution, and $p_{\theta}(x|z)$ is the Decoder output, reconstructing data from $z$.
While VAE has gained significant academic interest, its adoption in industry requires technological maturity and stability. It has been applied to anomaly detection in industrial settings [52, 53].
@@ II-C Generative Adversarial Network (GAN)
Generative Adversarial Networks (GANs) were first introduced by Goodfellow et al. [15] in 2014 as an unsupervised learning algorithm based on a zero-sum game between two neural networks: a generator (G) and a discriminator (D). Unlike traditional methods, GANs do not require prior knowledge or assumptions about the data distribution. Given real data $x$ and random noise $z$ sampled from prior distribution $p_{z}$, the objective function can be expressed as:
@@
Within industrial applications, GANs are widely used over the years [54, 55]. This progression underscores the increasing versatility and significance of GANs in enhancing industrial processes through innovative AI-driven solutions.
@@ II-D Diffusion Model
Diffusion models, inspired by the physical diffusion process, have gained prominence in machine learning. The Denoising Diffusion Probabilistic Model (DDPM) [19] has attracted significant academic attention, following the earlier score-based diffusion model proposed by Song et al. [56], later unified in SDE studies [57]. These models iteratively transform data into a simple distribution (e.g., Gaussian) and then reconstruct it via a two-stage Markov process.
For a given data distribution $p(x)$ , the forward diffusion process perturbs data over time as:
@@
where $\epsilon_{t}\sim\mathcal{N}(0,I)$ and $\beta_{t}$ is a predefined noise factor. A denoising function $q_{\theta}$ reconstructs $x_{t-1}$ from $x_{t}$ , leading to the DDPM loss function:
@@
Originally emerging in computer vision, diffusion models have expanded into sequential modeling [58], audio processing [59], and AI for scientific applications [60]. In industry, they have been applied to sensor data imputation [61] and anomaly detection [27], addressing complex real-world challenges.
@@ IV The proposed DGM-based AIGC framework and the current state-of-the-art DGMs
The DGM-baed AIGC framework is proposed for industrial data generation, as illustrated in Fig. 3. First, generative models learns the latent representations of industrial time series by modeling dynamic processes and correlations between variables, which have been introduced in Section II and Section III. Second, quality enhancement, sample augmentation, and privacy protection are performed to improve the quality and quantity of data, which will be discussed in Section IV. Finally, the generated samples will be applied to various industrial scenarios to address challenges such as sample scarcity in industrial scenarios, as presented in Section V. With the development of generative intelligence, there are many methods are proposed to tackle these issues.
@@ IV-A Limited Sample Augmentation
Limited sample augmentation utilizes generative methods to enhance datasets where sample sizes are insufficient for effective model training. This approach is critical in scenarios where acquiring real-world data is challenging, enabling the creation of realistic and diverse samples to improve model accuracy and generalization.
In industrial applications, this technique has significantly improved fault diagnosis and anomaly detection. A bidirectional alignment network with VAE addresses data scarcity in thermal power plant fault diagnosis, achieving generalization across rare fault categories [62]. MSGAN combines offline training with real-time inference, using an adaptive update strategy and gradient penalty to generate high-quality anomaly samples, enhancing anomaly detection in robotic sensors [63]. Recently, optimized diffusion models leveraging asymptotic denoising and temporal dynamics extraction [64, 65, 66] have outperformed GANs in predictive maintenance, further demonstrating their versatility in limited sample augmentation.
@@ IV-B Imbalanced Data Synthesis
Imbalanced data synthesis involves generating artificial data to address class imbalances in datasets, a crucial technique in machine learning and industrial applications, especially for underrepresented categories.
In industry, several innovative approaches tackle data imbalance. Fan et al. [35] proposed a VAE-based method for semiconductor fault detection, effectively mitigating the scarcity of defective wafer samples. DRL-GAN integrates reinforcement learning with GANs, achieving balanced data distributions [67]. Recently, diffusion models have gained traction in fault diagnosis. Yan et al. [20] introduced a noise-to-signal inversion mechanism for bearing fault detection, preserving spectral characteristics. PSDM enhances feature consistency in sparse fault patterns through adaptive weight sharing [68]. Additionally, digital twin architectures combined with diffusion processes embed physical constraints via multi-physics simulations, improving rotating machinery diagnostics [69].
@@ IV-C Sensor Signal Imputation
Several advanced methods have been developed for sensor data imputation. Velasco-Gallego et al. [28] introduced a VAE-based method for marine machinery sensor data, while FedTMI [70] applies edge-computed GANs and federated transfer learning to industrial missing data. Among GAN-based approaches, models like IM-GAN [71] and SGAIN [72] tackle complex multivariate time series and structural health monitoring, while Semi-GAN [73] and ST-GAIN [74] improve imputation in semiconductor and manufacturing data. DiffAD [27] integrates denoising diffusion with a multi-scale state space model for time series anomaly detection, capturing long-term dependencies. GD-GRU [75] combines a Gaussian diffusion process with GRU, enhancing environmental data repair.
@@ IV-D Sensor Signal Denoising
Sensor Signal Denoising has been explored through various models, including AR, GAN, and hybrid approaches. AR models have demonstrated effectiveness in denoising, with Liu et al. [29] integrating wavelet regional correlation thresholding and ARMA models to enhance acoustic emission signal clarity in industrial robots. In GAN-based methods, CycleGAN has been applied with Bayesian nonparametric estimation for online sensor signal denoising in milling processes [30]. Additionally, Att-DCDN leverages an encoder-decoder GAN structure to integrate synthetic and field seismic data, improving seismic noise attenuation. Min et al. [76] further enhance seismic data denoising by combining a pre-trained deep denoising autoencoder with transfer learning, optimizing training on limited field data.
@@
ATT-TSAAE.[85] 2024 Anomaly Detection GAN+Transformer
Time-Transformer[86] 2024 General Industrial Applications Transformer
DiffUCD[64] 2025 Fault Diagnosis Diffusion Models
SDE Diffusion [65] 2024 Predictive Maintenance Diffusion Models
Diff-MTS[66] 2024 Predictive Maintenance Diffusion Models
@@
Yang et al.[93] 2024 Fault Diagnosis Diffusion Models
PSDM[68] 2024 Fault Diagnosis Diffusion Models
Jiang et al.[69] 2024 Fault Diagnosis Diffusion Models
MBAC-GAN[94] 2025 Fault Diagnosis GAN
@@
DiffAD[27] 2023 Anomaly Detection Diffusion Models
@@
Sensor Signal Denoising Att-DCDN.[98] 2021 General Industrial Applications GAN